You are conducting a study to see if the proportion of women over 40 who regularly have mammograms is significantly less than 0.16. You use a significance level of
α=0.005
H0: p = 0.16
H1: p < 0.16
You obtain a sample of size
n = 458
in which there are 64 successes.
What is the test statistic for this sample? 3 decimal places
test statistic =
What is the p-value for this sample? 4 decimal places
p-value =

Answer by Boreal(15235)

p hat is 64/458=0.14
alpha is 0.05 (suppose) p{reject Ho|Ho true}
test statistic is z=(phat-po)/sqrt(p*(1-p))/n
this is -0.02/sqrt(0.16*0.84/458)
=-1.18 (not rounding p hat)
fail to reject Ho with p-value of 0.14 (not rounding)
if 0.14 is used for p hat, the values are -1.17 and 0.12